2DIR SPECTROSCOPY OF PERIPHERAL MEMBRANE PEPTIDE

\underline{ANN MARIE WOYS}, MARTIN ZANNI, 1101 University Ave., Madison, WI 53706.

2DIR spectroscopy is a good tool for studying peptide structure. Linewidth of transmembrane helices has been shown to vary with depth in the membrane. We suspect that 2D lineshapes should also be sensitive to different peripheral membrane peptides. We have synthesized the antibiotic, ovispirin, which is know from NMR to lie on the membrane surface. Results will be presented on 2D lineshape that address the orientation and depth of the helix as seen by amide I vibrational dynamics.